# Ordered random walks and the Airy line ensemble

### v1

@@ 1 Introduction
The existing results which are closest to the present paper are universality properties of last passage percolation in ‘thin’ rectangles considered in Bodineau and Martin [6] and Baik and Suidan [4]; and universality properties of ordered random walk bridges in Baik and Suidan [5]. Define $T(n,k)$ to be the last passage percolation time at the point $(n,k)\in\mathbb{Z}_{+}^{2}$ constructed from a random environment made up of i.i.d. random variables for a large class of distributions. Then [4, 6] show that $T(n,\lfloor n^{a}\rfloor)$ converges in a suitable scaling limit to the Tracy Widom GUE distribution whenever $a<3/7$. The main technique in the proof is an application of the Komlós-Major-Tusnády (KMT) coupling of random walks and Brownian motions. Baik and Suidan applied the same technique to study a slowly growing number $d$ of non-intersecting random walk bridges for a general class of increment distributions. The bridge condition is imposed at time 00 and $2$, there are $N$ random walk steps and the positions are evaluated at time $1$. They showed convergence of the top particle in a scaling limit to the Tracy Widom GUE distribution. The application of the KMT coupling is less accurate in this setting due to the need to account for the small spacings between particles. As a result, the convergence in [5] requires the condition $N\geq(2d)^{8d^{3}+4d^{2}}$.
@@ 2.1 Background and notation for ordered random walks
Let $(e_{ij})_{i\geq 1,1\leq j\leq d}$ be a collection of i.i.d. exponential random variables with rate $1$ and let $(X_{ij})_{i\geq 1,1\leq j\leq d}$ be a collection of i.i.d. random variables. For each $j=1,\ldots,d$ and $t\geq 0$ let
@@
Let $x=(x_{1},\ldots,x_{d})$ and $(S(t))_{t\geq 0}=(S_{1}(t),\ldots,S_{d}(t))_{t\geq 0}$. Let $W^{d}=\{(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}:x_{1}<x_{2}<\ldots<x_{d}\}$ denote the Weyl chamber.
Define the stopping time
@@
is the Vandermonde determinant. A direct computation shows that $V$ satisfies the harmonic condition and the main difficulty is to show that $V$ is finite and $V(x)>0$ for all $x\in W^{d}$. This was shown in [17, 23] in the discrete-time case; the continuous-time setting here can be proved in a similar way. Alternatively, we can use the methods developed in this paper: $V$ is integrable by Proposition 8 and $V(x)\sim\Delta(x)$ as $x_{j+1}-x_{j}\rightarrow\infty$ by Proposition 3 which is sufficient to show that $V(x)>0$ for all $x\in W^{d}$, see the proof of Proposition 4e in [17]. Once these conditions have been established $V$ can be used to construct a process in $W^{d}$ by a Doob $h$-transform. We refer to this process as an ordered random walk and let $\mathbf{E}^{(V)}$ denote the expectation under this change of measure: for bounded measurable $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$
@@
$\mathbf{E}_{x}^{(\Delta)}[\phi(B(t))]=\mathbf{E}_{x}\left[\phi(B(t))\frac{\Delta(B(t))}{\Delta(x)};\tau>t\right].$
@@ 2.3 Convergence of ordered random walks to the Airy line ensemble
Let $D^{N}(k,A)$ denote the set of the top $k$ lines at times in $A$. For any $k\geq 1$ and $L>0,$ Corwin and Hammond [10] showed that the line ensemble $D^{N}(k,[-L,L])$ converges weakly as $N\rightarrow\infty$ to $\mathcal{L}$ restricted to $\{1,\ldots,k\}\times[-L,L].$ This was proved for non-intersecting Brownian bridges in Theorem 3.1 of [10] (note the definition of the parabolic Airy line ensemble differs by a factor of $2^{1/2}$). This can be converted into a statement about non-intersecting Brownian motions by the usual transformations between Brownian bridges and Brownian motions. The convergence in the form stated here can be found in Theorem 2.1 of [14]. Note that the Airy line ensemble is absolutely continuous with respect to Brownian motions with diffusivity parameter $2$ meaning that the quadratic variation grows as $2(t-s)$ in an interval $[s,t].$ The original Brownian motions have diffusivity parameter $1$ but this is transformed to remove factors of $2^{1/2}$ in the scaling limit.
[…]
@@
The conventions in the existing literature are that the ordered random walks are ordered as $S_{1}(t)\leq\ldots\leq S_{d}(t)$ while in the Airy line ensemble $\mathcal{L}_{1}(t)>\mathcal{L}_{2}(t)>\ldots.$ We stick with both of these conventions at the expense of needing to define $X$ in terms of $S$ in a reversed order. Let $X^{T}(k,A)$ denote the set of the top $k$ lines at times in $A$. Our first result is the following.
@@
Suppose the conditions in Section 2.2 hold. Uniformly in $x\in W_{T,\varepsilon}$ and $d\leq T^{a}$ with $2a<1/2-\varepsilon$,
@@
The asymptotics in Proposition 3 could be combined with [8, 9, 38] in order to approximate ordered random walks by non-intersecting Brownian motions for a more general class of initial conditions including cases where the macroscopic limit is affected by the initial condition. We do not pursue this here since our focus is on universality with respect to the law of the random walks where the most challenging case is when all particles are started from zero. If we have $x\in W_{T,\varepsilon}$ in addition to the conditions for Theorem 1 then we can improve the value of $a$. The second part of the proof of Theorem 1 and thus the conditions for Lemma 18 are not needed. If we wanted to relax the condition $\lvert x_{j}\rvert=o(T^{1/2-a/6})$ then we could combine the approximation arguments used here with the results of [8, 9, 38]. Note that [8, 38] apply only at the level of marginal distributions rather than at the process level. The results of [9] apply to finite dimensional distributions which can then be combined with characterisations of the Airy line ensemble in terms of its finite dimensional distributions eg. [21].
@@ 2.4 Law of large numbers and fluctuations of linear statistics
We now consider macroscopic properties of the system of ordered random walks. In particular, we show that the law of large numbers and fluctuations of linear statistics agree with the case of non-intersecting Brownian motions. Let $\delta>0$ and define for $f:D[\delta,1]\to\mathbb{R}$
@@
Suppose the conditions in Section 2.2 hold. Let $\delta>0$ and suppose that $f:D[\delta,1]\rightarrow\mathbb{R}$ is uniformly continuous and $g:\mathbb{R}\rightarrow[0,1]$ is uniformly continuous. For $d=\lfloor T^{a}\rfloor$ with $a<1/16$ and all starting positions $\lvert x_{j}\rvert=o(T^{1/2+a/2})$
@@
Suppose the conditions in Section 2.2 hold. Let $\delta>0$ and suppose that $f:D[\delta,1]\rightarrow\mathbb{R}$ is uniformly continuous and $g:\mathbb{R}\rightarrow[0,1]$ is uniformly continuous. For $d=\lfloor T^{a}\rfloor$ with $a<1/16$ and all starting positions $\lvert x_{j}\rvert=o(T^{1/2+a/2})$
@@
At a fixed time, the fluctuations of linear statistics of eigenvalues of random matrices have been considered in many papers starting with the work of [20, 28]. For processes evolving in time, the fluctuations of linear statistics have been studied by Duits [22] for non-intersecting Brownian bridges where the function $f:C[0,1]\rightarrow\mathbb{R}$ has the form $f(w)=\sum_{i=1}^{N}f_{i}(w(t_{i}))$ for $10<t_{1}<\ldots<t_{N}<$ 1 where $x\rightarrow f_{i}(x)$ for all $i=1,\ldots,N$ is continuously differentiable and grows at most polynomially for $x\rightarrow\pm\infty$. After applying time changes to the setting of non-intersecting Brownian motions, the results of [22] show for this class of functions that
@@ 3 Superharmonic functions
We can replace Condition (ii) in Section 2.2 with the following condition that we call Condition (ii)’. Suppose that we can find a random variable $\zeta$ taking values in $[1,\infty)$ such that for all $\theta\geq 0$,
@@
Then $\zeta$ takes values in $[1,\infty)$ and satisfies (4) and (5).
Let $\eta_{1}:=0$ and for $j=2,\ldots,d$ let $\eta_{j}:=\sum_{i=1}^{j-1}\zeta_{i}$ where $(\zeta_{i})_{i=1}^{d-1}$ are i.i.d. with common distribution $\zeta$. Let $\eta=(\eta_{1},\ldots,\eta_{d})\in W^{d}$.
@@
Suppose Condition (ii)’ holds. For $x\in W^{d},$ let
@@
Suppose that $U,W$ are independent positive random variables such that $U\stackrel{{\scriptstyle lr}}{{\leq}}W$. Let $\phi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ satisfy (i) $\phi(u,w)+\phi(w,u)\geq 0$ for all $w\geq u\geq 0$ and (ii) $\phi(u,w)\geq 0$ for all $w\geq u\geq 0$. Then
@@
Suppose that we can show that for $k=1,\ldots,d-1$,
@@
Here recall that only one walker can jump at a time. Using Theorem 2.1 from [31] that $\Delta$ is harmonic for $(S(t))_{t\geq 0}$ (here there is no killing) then $h$ is also harmonic for $(S(t))_{t\geq 0}$ since
$\mathbf{E}_{x}[h(S(t)]$ $=\mathbf{E}_{x}\mathbf{E}^{\eta}[\Delta(S(t)+x+\eta)]=\mathbf{E}^{\eta}\mathbf{E}_{x}[\Delta(S(t)+x+\eta)]$
@@
$h(x)$ $\geq\limsup_{n\rightarrow\infty}\mathbf{E}_{x}[h(S(t))1_{\{\tau>t\}}]$
$\geq\limsup_{n\rightarrow\infty}\mathbf{E}_{x}[\Delta(S(t))1_{\{\tau>t\}}]$
$\geq V(x).$
As a result it is sufficient to prove (6) and (7). For $k=1,\ldots,d-1$ let $Y=x_{k}+X_{k}-x_{k+1}|(x_{k}+X_{k}>x_{k+1})$ and $Z=\zeta_{k}$. Consider
@@
where $c_{p,q}\geq 0$ is a constant and $\psi:\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ is symmetric and positive. For any $p,q\in\mathbb{Z}_{\geq 0}$ we have that $\phi_{p,q}(y,z)\geq 0$ for all $z\geq y\geq 0$ as each factor is positive. Moreover,

### v2

# arXiv:2411.17827v2[math.PR] 03 Mar 2026

@@ 1 Introduction
The existing results which are closest to the present paper are universality properties of last passage percolation in ‘thin’ rectangles considered in Bodineau and Martin [9] and Baik and Suidan [5]; and universality properties of ordered random walk bridges in Baik and Suidan [6]. Define $T(n,k)$ to be the last passage percolation time at the point $(n,k)\in\mathbb{Z}_{+}^{2}$ constructed from a random environment made up of i.i.d. random variables for a large class of distributions. Then [5, 9] show that $T(n,\lfloor n^{a}\rfloor)$ converges in a suitable scaling limit to the Tracy Widom GUE distribution whenever $a<3/7$. The main technique in the proof is an application of the Komlós-Major-Tusnády (KMT) coupling of random walks and Brownian motions. Baik and Suidan applied the same technique to study a slowly growing number $d$ of non-intersecting random walk bridges for a general class of increment distributions. The bridge condition is imposed at time 0 and $2$, there are $N$ random walk steps and the positions are evaluated at time $1$. They showed convergence of the top particle in a scaling limit to the Tracy Widom GUE distribution. The application of the KMT coupling is less accurate in this setting due to the need to account for the small spacings between particles. As a result, the convergence in [6] requires the condition $N\geq(2d)^{8d^{3}+4d^{2}}$.
@@ 2.1 Background and notation for ordered random walks
Let $(e_{ij})_{i\geq 1,1\leq j\leq d}$ be a collection of i.i.d. exponential random variables with rate $1$ and let $(X_{ij})_{i\geq 1,1\leq j\leq d}$ be a collection of i.i.d. random variables. Let $x=(x_{1},\ldots,x_{d})$ and $(S(t))_{t\geq 0}=(S_{1}(t),\ldots,S_{d}(t))_{t\geq 0}$. For each $j=1,\ldots,d$ and $t\geq 0$ let
@@
Let $W^{d}=\{(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}:x_{1}<x_{2}<\ldots<x_{d}\}$ denote the Weyl chamber. Define the stopping time
@@
is the Vandermonde determinant. A direct computation shows that $V$ satisfies the harmonic condition and the main difficulty is to show that $V$ is finite and $V(x)>0$ for all $x\in W^{d}$. This was shown in [21, 26] in the discrete-time case; the continuous-time setting here can be proved in a similar way. Alternatively, we can use the methods developed in this paper: $V$ is integrable by Proposition 8 and $V(x)\sim\Delta(x)$ as $x_{j+1}-x_{j}\rightarrow\infty$ by Proposition 3 which is sufficient to show that $V(x)>0$ for all $x\in W^{d}$, see the proof of Proposition 4e in [21]. Once these conditions have been established $V$ can be used to construct a process in $W^{d}$ by a Doob $h$-transform. We refer to this process as an ordered random walk and let $\mathbf{E}^{(V)}$ denote the expectation under this change of measure: for bounded measurable $\phi:\mathbb{R}^{n}\rightarrow,\mathbb{R}$
@@
$\mathbf{E}_{x}^{(\Delta)}[\phi(B(t))]=\mathbf{E}_{x}\left[\phi(B(t))\frac{\Delta(B(t))}{\Delta(x)};\tau^{\text{bm}}>t\right].$
@@ 2.3 Convergence of ordered random walks to the Airy line ensemble
Let $D^{N}(k,A)$ denote the set of the top $k$ lines at times in $A$. For any $k\geq 1$ and $L>0,$ Corwin and Hammond [13] showed that the line ensemble $D^{N}(k,[-L,L])$ converges weakly as $N\rightarrow\infty$ to $\mathcal{L}$ restricted to $\{1,\ldots,k\}\times[-L,L].$ This was proved for non-intersecting Brownian bridges in Theorem 3.1 of [13] (note the definition of the parabolic Airy line ensemble differs by a factor of $2^{1/2}$). This can be converted into a statement about non-intersecting Brownian motions by the usual transformations between Brownian bridges and Brownian motions. The convergence in the form stated here can be found in Theorem 2.1 of [17]. Note that the Airy line ensemble is absolutely continuous with respect to Brownian motions with diffusivity parameter $2$ meaning that the quadratic variation grows as 2(ts)2(t-s) in an interval $[s,t].$ The original Brownian motions have diffusivity parameter $1$ but this is transformed to remove factors of $2^{1/2}$ in the scaling limit.
Furthermore let $N=\lfloor T^{a}\rfloor$ (we omit $\lfloor\cdot\rfloor$ to simplify the notation) and use Brownian scaling to rewrite (2) as
@@
The conventions in the existing literature are that the ordered random walks are ordered as $S_{1}(t)\leq\ldots\leq S_{d}(t)$ while in the Airy line ensemble $\mathcal{L}_{1}(t)>\mathcal{L}_{2}(t)>\ldots.$ We stick with both of these conventions at the expense of needing to define $X$ in terms of $S$ in a reversed order.
Let $X^{T}(k,A)$ denote the top $k$ lines at times in $A\subset\mathbb{R}$ as an element of $D(k,A)$, the set of $k$-tuples of càdlàg functions on $A$ with the $\sigma$-algebra generated by the finite-dimensional evaluation maps. We consider the uniform metric on $D(k,A)$,
$\rho(x,y)=\sup_{1\leq j\leq k}\sup_{t\in A}\lvert x_{k}(t)-y_{k}(t)\rvert$ (4)
for $x=(x_{1}(t),\ldots,x_{k}(t))_{t\in A},y=(y_{1}(t),\ldots,y_{k}(t))_{t\in A}\in D(k,A)$. We consider weak convergence in this setting, in the sense of Section 6 and Section 15 of [8]. Alternatively, we could restrict $X^{T}(k,A)$ by taking a linear interpolation to the set $C(k,A)$ of $k$-tuples of continuous functions on $A$, with the uniform topology, and weak convergence would follow in this setting after some adaptions to Proposition 15.
Our first result is the following.
@@
Suppose the conditions in Section 2.2 hold, $a\in(0,1/4)$ and $\varepsilon>0$. Uniformly in $x\in W_{T,\varepsilon}$ and $d\leq T^{a}$ with $2a<1/2-\varepsilon$,
@@
Our focus is on universality with respect to the law of the random walks where the most challenging case is when all particles are started from zero. The arguments we give to approximate functionals of ordered random walks by non-intersecting Brownian motions hold more generally, including when the limiting object is not the Airy line ensemble. In Section 2.4 we state two results involving macroscopic limits. Equally our methods could also be applied to other microscopic limits such as the Dyson sine process, non-intersecting Brownian motions with outliers or the Pearcey process. It is worth noting that the convergence of non-intersecting Brownian motions to a limiting object such as the Airy line ensemble is only known in particular cases, e.g. for some special choices of initial conditions or in the sense of finite-dimensional distributions.
The asymptotics in Proposition 3 could also be combined with papers [11, 12, 41] that prove convergence in the edge scaling limit of the one-point or finite-dimensional distributions of non-intersecting Brownian motions to the Airy line ensemble for a more general class of initial conditions. If we have $x\in W_{T,\varepsilon}$ in addition to the conditions for Theorem 1 then we can improve the value of $a$. The second part of the proof of Theorem 1 and thus the conditions for Lemma 18 are not needed. If we wanted to relax the condition $\lvert x_{j}\rvert=o(T^{1/2-a/6})$ then we could combine the approximation arguments used here with the results of [11, 12, 41]. Note that [11, 41] apply only at the level of marginal distributions rather than at the process level. The results of [12] apply to finite-dimensional distributions which can then be combined with characterisations of the Airy line ensemble in terms of its finite dimensional distributions, e.g. [24].
@@ 2.4 Law of large numbers and fluctuations of linear statistics
We now consider macroscopic properties of the system of ordered random walks. In particular, we show that the law of large numbers and fluctuations of linear statistics agree with the case of non-intersecting Brownian motions. Let $D([\delta,1])$ denote the set of càdlàg functions with the sigma-algebra generated by the finite-dimensional evaluation maps and the uniform metric $\rho$ given in (4). Let $\delta>0$ and define for $f:D([\delta,1])\to\mathbb{[}0,1]$,
@@
Suppose the conditions in Section 2.2 hold. Let $\delta>0$ and suppose that $f:D([\delta,1])\to\mathbb{R}$ is measurable, uniformly continuous and $g:\mathbb{R}\rightarrow[0,1]$ is uniformly continuous. For $d=\lfloor T^{a}\rfloor$ with $a<1/16$ and all starting positions $\lvert x_{j}\rvert=o(T^{1/2+a/2})$,
@@
Suppose the conditions in Section 2.2 hold. Let $\delta>0$ and suppose that $f:D[\delta,1]\rightarrow\mathbb{R}$ is measurable, uniformly continuous and $g:\mathbb{R}\rightarrow[0,1]$ is uniformly continuous. For $d=\lfloor T^{a}\rfloor$ with $a<1/16$ and all starting positions $\lvert x_{j}\rvert=o(T^{1/2+a/2})$,
@@
At a fixed time, the fluctuations of linear statistics of eigenvalues of random matrices have been considered in many papers starting with the work of [23, 31]. For processes evolving in time, the fluctuations of linear statistics have been studied by Duits [25] for non-intersecting Brownian bridges where the function $f:C[0,1]\rightarrow\mathbb{R}$ has the form $f(w)=\sum_{i=1}^{N}f_{i}(w(t_{i}))$ for $0<t_{1}<\ldots<t_{N}<1$ where $x\rightarrow f_{i}(x)$ for all $i=1,\ldots,N$ is continuously differentiable and grows at most polynomially for $x\rightarrow\pm\infty$. After applying time changes to the setting of non-intersecting Brownian motions, the results of [25] show for this class of functions that
@@ 3 Superharmonic functions
We can replace Condition (ii) in Section 2.2 with the following condition that we call Condition (ii)’. Let $X$ denote a random variable with the same law as $(X_{ij})_{i\geq 1,1\leq j\leq d}$. Let $\zeta$ be a random variable taking values in $1,\infty)$ such that there exists $\delta_{0}>0$ such that $\mathbf{E}[e^{\delta\zeta}]<\infty$ for all $\delta<\delta_{0}.$ If there exists $M>0$ such that $\mathbf{P}(\lvert X\rvert\geq M)=0$ then we set $\zeta:=M$. Otherwise, suppose that we can find $\zeta$ satisfying such that for all $\theta\geq 0$,
@@
Here, $(X-\theta)|\{X>\theta\}$ is the random variable whose law is that of $X-\theta$, conditioned on the event $X>\theta$. If $X$ is almost surely bounded above (resp. below) by M then we take $\zeta\geq M$ satisfying (6) (resp. (5)).
@@
Here, (X|X>0)(X|X<0)(X|X>0)-(X|X<0) is denoting the difference of two independent random variables, whose laws are $X$ conditioned on $X>0$ and $X<0$ respectively. Then, $\zeta$ takes values in $1,\infty),$ has a finite exponential moment and satisfies (5) and (6).
Let $\eta_{1}:=0$ and for $j=2,\ldots,d$ let $\eta_{j}:=\sum_{i=1}^{j-1}\zeta_{i}$ where $(\zeta_{i})_{i=1}^{d-1}$ are i.i.d. with common distribution $\zeta$ and are also independent of any of the $(X_{ij})_{i\geq 1,1\leq j\leq d}$. Let $\eta=(\eta_{1},\ldots,\eta_{d})\in W^{d}$.
@@
Suppose the conditions in Section 2.2 hold, where either Condition (ii) or (ii)’ can be imposed. For $x\in W^{d},$ let
@@
Suppose that $U,W$ are independent positive random variables such that $U\stackrel{{\scriptstyle lr}}{{\leq}}W$. Let $\phi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ satisfy (i) $\phi(u,w)+\phi(w,u)\geq 0$ for all $w\geq u\geq 0$ and (ii) $\phi(u,w)\geq 0$ for all $w\geq u\geq 0$. Then,
@@
Let $X_{k}$ denote a random variable with the same law as $(X_{ij})_{i\geq 1,1\leq j\leq d}.$ Suppose that we can show that for $k=1,\ldots,d-1$,
@@
Here, recall that only one walker can jump at a time. Using Theorem 2.1 from [33] that $\Delta$ is harmonic for $(S(t))_{t\geq 0}$ (here there is no killing) then $h$ is also harmonic for $(S(t))_{t\geq 0}$ since
$\mathbf{E}_{x}[h(S(t))]$ $=\mathbf{E}_{x}\mathbf{E}^{\eta}[\Delta(S(t)+\eta)]=\mathbf{E}^{\eta}\mathbf{E}_{x}[\Delta(S(t)+\eta)]$
@@
$h(x)$ $\geq\limsup_{t\rightarrow\infty}\mathbf{E}_{x}[h(S(t))1_{\{\tau>t\}}]$
$\geq\limsup_{t\rightarrow\infty}\mathbf{E}_{x}[\Delta(S(t))1_{\{\tau>t\}}]$
$=V(x).$
As a result, it is sufficient to prove (7) and (8). If $X_{k}$ is almost surely bounded then the equations are automatically satisfied. For $k=1,\ldots,d-1$ let $Y=x_{k}+X_{k}-x_{k+1}|(x_{k}+X_{k}>x_{k+1})$ and $Z=\zeta_{k}$. Note that $h$ is an expectation of a product indexed by $(i,j)$ for $1\leq i<j\leq d.$ We can write this product with the indices in the order $(k,k+1)$, $(j,k)$ for $j<k$, $(j,k+1)$ for $j<k+1$, $(k,j)$ for $j>k+1$, $(k+1,j)$ for $j>k+1$ and finally $(i,j)$ for $i<j\notin\{k,k+1\}.$ This gives
@@
where $c_{p,q}\geq 0$ is a constant and $\psi:\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ is symmetric and positive. The symmetric function $\psi$ arises from the symmetry between the factors in the product for $h$ indexed by $(j,k)$ for $j<k$ and $(j,k+1)$ for $j<k$. For any $p,q\in\mathbb{Z}_{\geq 0}$ we have that $\phi_{p,q}(y,z)\geq 0$ for all $z\geq y\geq 0$ as each factor is positive. Moreover,